OpenStudy (anonymous):

take lim x approaching -infinity (1-2x)/(sqrt(x^2+x)) for example, why would we only divide the bottom by x^2 and not the top?

OpenStudy (anonymous):

eyeball problem

OpenStudy (anonymous):

think of it as \[\frac{-2x}{x}\]

OpenStudy (anonymous):

if you want to do it the math teacher way, divide numerator and denominator by \(x\) but that is mostly a waste of time

OpenStudy (anonymous):

\[\frac{\frac{1}{x}-2}{\sqrt{1+\frac{1}{x}}}\]

OpenStudy (anonymous):

confused

OpenStudy (anonymous):

where?

OpenStudy (anonymous):

the 1 in the numerator is not important

OpenStudy (anonymous):

neither is the \(x\) in the square root in the denominator

OpenStudy (anonymous):

why not

OpenStudy (anonymous):

it is determined entirely by the highest degree in the top and bottom

OpenStudy (anonymous):

just like \[\lim_{x\to \infty}\frac{2x^2+3x-4}{3x^2+5x-20}=\frac{2}{3}\]

OpenStudy (anonymous):

i get that , but take this other one for example , lim x approaching infinity (x^2-3x+7)/(x^3+10x-4), you would divide the top and bottom by x^3, but in the one i showed you first, the top is divided by x?

OpenStudy (anonymous):

i thought the highest one in the denominator was the one used to divide

OpenStudy (anonymous):

i would do none of those things in the example you just wrote the degree of the denominator is larger than the degree of the numerator, making the horizontal asympote \(y=0\) bigger denominator means smaller number (small as in close to zero)

OpenStudy (anonymous):

if you want to divide by \(x^3\) you will get the same answer

OpenStudy (anonymous):

if it like x^2/x^3 and x is approaching infinity , it would be 0

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

ok, but in the first one, if you divide by x^2 on top and bottom, you wont get -2

OpenStudy (anonymous):

OOOHO i see the problem

OpenStudy (anonymous):

you have in the denomiantor \(\sqrt{x^2+x}\) not \(x^2+x\)

OpenStudy (anonymous):

then when you divide by \(x\) it comes inside the radical as a \(x^2\)

OpenStudy (anonymous):

i.e. \[\frac{\sqrt{x^2+x}}{x}=\sqrt{\frac{x^2}{x^2}+\frac{x}{x^2}}\]

OpenStudy (anonymous):

so when you divide top and bottom you get literally \[\frac{\frac{1}{x}-\frac{2x}{x}}{\sqrt{1+\frac{1}{x}}}\]

OpenStudy (anonymous):

ok so this only works when there is a square root in either bottom or top, you can divide it by lets say the top x and the bottom x^2?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

you divide top and bottom by \(x\)

OpenStudy (anonymous):

but as is said, the x comes inside the radical as a \(x^2\)

OpenStudy (anonymous):

\[\frac{\sqrt{x^2+x}}{x}=\sqrt{\frac{x^2}{x^2}+\frac{x}{x^2}}\]

OpenStudy (anonymous):

ok, one last thing, lets change the top x with x^3 , what would we divide it by

OpenStudy (anonymous):

if the numerator was \(x^3\) then the limit would be infinite, since the degree of the top would be larger than the degree of the bottom

OpenStudy (anonymous):

btw your answer is \(2\) since you are taking the limit as x goes to \(-\infty\) not \(\infty\)

OpenStudy (anonymous):

oh okay thanks

OpenStudy (anonymous):

yw